Find the inverse Laplace transform of In s+3/s^2+9?


1
Expert's answer
2021-09-03T01:59:56-0400

"L^{-1}\left\{\frac{s+3}{s^2+9}\right\}\\
\frac{s+3}{s^2+9}=\frac{s}{s^2+9}+\frac{3}{s^2+9}\\
=L^{-1}\left\{\frac{s}{s^2+9}+\frac{3}{s^2+9}\right\}\\
\mathrm{Use\:the\:linearity\:property\:of\:Inverse\:Laplace\:Transform:}\\
\mathrm{For\:functions\:}f\left(s\right),\:g\left(s\right)\mathrm{\:and\:constants\:}a,\:b:\quad L^{-1}\left\{a\cdot f\left(s\right)+b\cdot g\left(s\right)\right\}=a\cdot L^{-1}\left\{f\left(s\right)\right\}+b\cdot L^{-1}\left\{g\left(s\right)\right\}\\
=L^{-1}\left\{\frac{s}{s^2+9}\right\}+3L^{-1}\left\{\frac{1}{s^2+9}\right\}=\cos \left(3t\right)+3\cdot \\\frac{1}{3}\sin \left(3t\right)\\
=\cos \left(3t\right)+\sin \left(3t\right)"